OpenStudy (anonymous):

Last one :P Mia deposits either $5 or $3.50 into her savings account. She made 14 deposits equaling a total of $59.50. How many times did she deposit $5 and how many times did she deposit $3.50? Mia made deposits of $5 and deposits of $3.50.

OpenStudy (jdoe0001):

she made 14 deposits TOTALing 59.50 so whatever amount was for $5, let's say "x" so whatever amount was for $3.50, let's say "y" whatever those amounts are, they both sum up to 14, since it was only 14 deposits \(\Large \bf x +y =14\) how much was that in money? well "x" amount at $5 each is 5x what about the 3.50? well "y" at $3.50 is 3.50y whatever amount that was, TOTALed 59.50 thus \(\Large \bf 5x + 3.50y = 59.50\)

OpenStudy (ilovecake):

You need to make a system of equations :) x = number of times she deposited 5 dollars y = number of times she deposited $3.50 Equation One: 5x + 3.5y = 59.5 Equation Two: x + y = 14 I'm using the substitution method, because I find that to be the fastest. For equation two, subtract y from both sides ( (x+y)-y = 14 - y) You'll get x = 14 - y Substitute the new equation two in for x in equation one: 5(14-y) + 3.5y = 59.5 70 - 5y + 3.5y = 59.5 70 - 1.5y = 59.5 -1.5y = -10.5 -1.5y/-1.5 = -10.5/-1.5 y = 7 Plug in y for the original equation two, x + 7 = 14 x = 7 Mia deposited $5 seven times and $3.50 seven times.

OpenStudy (ilovecake):

In other words: x number of deposits for $5 and y ......................................... for $3.5 total deposits x+y=14 5x number of deposit for $5 3.5y ....................... for $3.50 then the second equation is 5x+3.5y=59.50

OpenStudy (ilovecake):

oh you gave the other guy a medal. I wasted my time! :(